# LLMInit: A Free Lunch from Large Language Models for Selective Initialization of Recommendation

### v1

@@ 4.3. Efficiency Comparison of LLM-based RecSys
Table 3 compares various LLM-based RecSys in full-ranking settings, focusing on the computation cost of parameter size and performances on the Beauty and Toys datasets. For the LLMRank (Hou et al., 2024) and LLMRec (Wei et al., 2024), we adopt the LLaMa-7B  [^3] https://huggingface.co/meta-llama/Llama-2-7b model for text generation. MoRec(Yuan et al., 2023), a modality-based model, leverages LLMs to generate textual embeddings. However, its performance is hindered by the use of a unified transform layer for embedding mapping to each item. Additionally, the parameter size for LLM-generated embeddings, initially around 13M/14M, can scale to billions as the total number of users and items surpasses 10 million, posing significant computational and storage challenges. Note that, LLMRank (Hou et al., 2024) directly employs an LLM for prompting but struggles with an incomplete long-context candidate pool and hallucination problem, leading to unreported performance results. LLMRec incorporates LLM for graph augmentation, improving recall to 0.0974 (Beauty) and 0.082 (Toys) but at a much higher computational cost (7B+ parameters). LLMInit, a lightweight alternative, achieves the best performance with only 0.2M parameters, highlighting the potential of parameter-efficient strategies in LLM-powered RecSys.
4.4. Cold-Start Recommendation
@@
Figure 3. Performance improvement comparison in non-strict cold-start scenarios on Amazon-Beauty, where half of the observed interactions for each user are removed.
In the long-tail cold-start setting (Zhang et al., 2025), we remove half of the observed interactions of each user in training, resulting in many cold-start users with only a single interaction. Figure 3 highlights how the three CF models—LightGCN, SGL, and SGCL— react to the limited interactions and benefit from LLM-based initialization strategies in cold-start scenarios. LightGCN, as a simple graph model, is most resistant to the cold-start challenge, while the SGL and SGCL both suffer sharp performance drops due to reliance on interactions for data augmentation or supervised loss optimization. With the LLMInit, SGL demonstrates a stronger ability to utilize enhanced embeddings, by extending LightGCN with self-supervised loss. SGCL, which introduces supervised contrastive learning loss, achieves the highest gains with LLMInit strategies, with Recall and NDCG improving by over 50.4% and 58.9%, respectively. The supervised loss in SGCL amplifies the impact of informative embeddings, allowing it to extract more meaningful patterns from the selected dimensions. Overall, the results reveal that more advanced models like SGCL with nuanced supervised graph contrastive loss are better equipped to exploit high-quality LLM-based initializations.
4.5. The impacts of Large Language Models
@@
Table 4. The Performance of LightGCN on Amazon-Beauty with initialization of different LLMs.
LLMs MPNet4441 gte-Qwen5552 GPT-S6663 GPT-L7773 Stella8884
@@
Rand R@10 0.0960 0.0936 0.0967 0.0925 0.0996
N@10 0.0467 0.0418 0.0463 0.044 0.0447
Uni R@10 0.1006 0.096 0.098 0.0942 0.0949
N@10 0.0469 0.0428 0.0465 0.0445 0.0417
Var R@10 0.1019 0.0932 0.098 0.096 0.0935
N@10 0.0485 0.0429 0.0465 0.046 0.0412
Table 4 illustrates the performance of LightGCN with different LLMs and initialization strategies (Rand, Uni, Var) on Amazon-Beauty. It observed that all LLMs can enhance LightGCN significantly with proper initialization strategies. Variance-based initialization (Var) generally outperforms random (Rand) and uniform (Uni) strategies by prioritizing high-variance dimensions, capturing the most informative features. Notably, MPNet, with the smallest embedding size (768), achieves the highest Recall@10 (0.1019) and NDCG@10 (0.0485), highlighting the importance of embedding quality over size. While larger models like GPT-L and Stella improve LightGCN with LLMInit, they still lag behind MPNet and GPT-S, emphasizing that LLM alignment with the dataset and strategic initialization matter more than embedding size alone.
5. Conclusion
In this work, we propose LLMInit, an effective and scalable framework that leverages pre-trained large language model embeddings for collaborative filtering model initialization. By employing selective strategies for item embeddings and aggregation pooling for user embeddings, LLMInit addresses data sparsity and cold-start challenges while ensuring scalability and efficiency with low-resource embedding size. Extensive experiments show LLMIint consistently enhancing diverse SOTA CF models, demonstrating LLMs’ new potential research direction to advance recommender systems.

### v2

@@ 4.3 Efficiency and Perforamance Comparisons of LLM-based RecSys
Table 2 compares computation cost and performance of various LLM-based RecSys in full-ranking settings. For the LLMRank hou2024large and LLMRec wei2024llmrec, we adopt the LLaMa-7B model for text generation. LLMRank hou2024large directly employs an LLM for prompting but struggles with an incomplete long-context candidate pool and hallucination problem, leading to unreported performance results. LLMRec prompts LLM for graph augmentation, achieving recall 0.0974 (Beauty) and 0.082 (Toys) with high computational cost (7B+ parameters). In comparison, MoRecyuan2023go, a modality-based model, leverages LLMs to generate textual embeddings but restricted by the use of a unified transform layer for distinct item embedding mapping. In addition, the parameter size for LLM-generated embeddings, initially around 13M/14M for MoRec and LLMInit-Full, can scale to billions as the total number of users and items surpasses 1 million, posing significant computational and storage challenges. TIGER, employing vector quantization for generative recommendation, shows potential for scalable RecSys. LLMInit-Var, a lightweight alternative, achieves the best performance with only 2M parameters, highlighting the potential of parameter-efficient strategies in LLM-based RecSys.
@@
Figure 3: Performance improvement comparison in non-strict cold-start scenarios on Amazon-Beauty, where half of the observed interactions for each user are removed.
@@
Table 3: The Performance of LightGCN on Amazon-Beauty with initialization of different LLMs. The embedding size of LLMs scale from 768 to 8,096 while LLMInit models all adopt the embedding size of 128, efficient for large user base scenarios.
LLMs MPNet2222 gte-Qwen3333 GPT-S4444 GPT-L5555 Stella6666
@@
LLMInit-Rand Recall@10 0.0960 0.0936 0.0967 0.0925 0.0996
NDCG@10 0.0467 0.0418 0.0463 0.044 0.0447
LLMInit-Uni Recall@10 0.1006 0.096 0.098 0.0942 0.0949
NDCG@10 0.0469 0.0428 0.0465 0.0445 0.0417
LLMInit-Var Recall@10 0.1019 0.0932 0.098 0.096 0.0935
NDCG@10 0.0485 0.0429 0.0465 0.046 0.0412
4.4 Cold-Start Recommendation
In the long-tail cold-start setting zhang2025cold; yang2025cold, we remove half of the observed interactions of each user in training, resulting in many cold-start users with only a single interaction. Figure 3 highlights how the three CF models (LightGCN, SGL, and SGCL) react to the limited interactions and benefit from LLM-based initialization strategies in cold-start scenarios. LightGCN, as a simple graph model, is most resistant to the cold-start challenge, while the SGL and SGCL both suffer sharp performance drops due to reliance on interactions for data augmentation or supervised loss optimization. With the LLMInit, SGL and SGCL demonstrate stronger abilities to utilize enhanced embeddings. In particular, SGCL achieves the highest gains with LLMInit strategies, with Recall and NDCG improving by over 50.4% and 58.9%, respectively. The supervised loss in SGCL amplifies the impact of informative embeddings, allowing it to extract more meaningful patterns from the selected dimensions. Overall, the results reveal that more advanced models like SGCL with nuanced supervised graph contrastive loss are better equipped to exploit high-quality LLM-based initializations.
4.5 The impacts of Large Language Models
Table 3 presents the performance of LightGCN initialized with embeddings from various LLMs using different strategies on the Amazon-Beauty dataset. The results show that all tested LLMs can enhance LightGCN’s performance when paired with appropriate initialization methods. Among them, the variance-based strategy consistently delivers superior results by selecting high-variance dimensions that retain the most informative semantic features. While larger models such as GPT-L and Stella provide higher-dimensional embeddings, their advantages are not linearly correlated with model size. Instead, MPNet and GPT-S exhibit superior performance, suggesting that the quality and domain alignment of embeddings play a more decisive role than sheer model capacity. Consequently, optimizing initialization and embedding relevance emerges as a more effective approach than simply increasing embedding dimensionality or model size for large-scale recommendation scenarios.
5 Conclusion
In this work, we introduced LLMInit, a practical and scalable framework designed for industrial LLM RecSys, which effectively integrates LLM embeddings for CF model initialization. By employing selective strategies for item embeddings and aggregation pooling for user embeddings, LLMInit addresses critical industry challenges of existing CF models, such as data sparsity and cold-start scenarios, while maintaining computational efficiency suitable for large-scale deployments. Extensive evaluations on real-world datasets validate that LLMInit significantly boosts the performance of various state-of-the-art CF models, demonstrating the tangible value and feasibility of leveraging language-based world knowledge to improve recommendation quality in industrial applications.
Acknowledgment
This work is supported in part by NSF under grants III-2106758, and POSE-2346158.
Limitations
Although LLMInit achieves notable improvements for scalable LLM-based recommendation, especially under cold-start and sparse-data scenarios, its effectiveness relies on the quality and domain relevance of pretrained LLM embeddings, necessitating careful selection of LLMs for domains or conducting further domain adaptation. Additionally, our current approach leverages only textual metadata; incorporating multi-modal features (such as images or audio) remains an important future direction to enhance real-world applicability. Finally, more advanced CF models such ordinary-differential-equations (ODEs) xu2023graph; xu2025graph; zhang2024we have not yet been evaluated within LLMInit.
Appendix A Dataset Statistics
Table 4: Statistics of the datasets.
Dataset # Users # Items # Inter. Sparsity Beauty 10,554 6,087 94,148 99.85% Toys–Games 11,269 7,310 95,420 99.88% Tools–Home 9,246 6,199 73,250 99.87% Office–Products 3,261 1,584 36,544 99.29%
To comprehensively evaluate the effectiveness of LLM-based initialization strategies in real-world recommendation scenarios, we conduct experiments on four widely used Amazon product review datasets (ni2019justifying), covering diverse user interaction domains. Each dataset consists of implicit user–item interactions, where an interaction denotes a user’s engagement (e.g., rating and review) with a product. Following common practice, we filter out users and items with fewer than five interactions and adopt the leave-one-out evaluation protocol. The statistics of the processed datasets are summarized in Table 4.
Appendix B Pooling Strategies
Table 5: Performance comparison of different initialization pooling strategies on Amazon-Beauty. All models are initialized with LLM-based embeddings except for the random baseline.
Method Recall@10 NDCG@10
Random Init 0.1029 0.0475
Mean (LLMInit) 0.1104 0.0522
Max (LLMInit) 0.1088 0.0510
LightGCN (LLMInit) 0.1101 0.0515
To further analyze the impact of different embedding aggregation strategies, Table 5 compares pooling-based methods for user embedding initializations. Results show that all LLM-based initialization strategies outperform random initialization, confirming the benefits of semantic priors learned by LLMs. Mean pooling achieves the best overall performance, slightly surpassing both max pooling and LightGCN pooling. This indicates that averaging item embeddings provides a stable and representative initialization signal for graph-based user–item interaction modeling.